OpenStudy (sasogeek):

integrate sex^2 xe^tanx dx from 0 to 1

OpenStudy (sasogeek):

\[\int\limits_{0}^{1}\sec ^{2}xe ^{tanx}dx\]

OpenStudy (sasogeek):

oops.... not "sex".... sec* ma bad...

OpenStudy (anonymous):

substitue tanx=t so u'll get sec^2xdx=dt and so u now need to integrate just e^t from 0 to pi/4.

OpenStudy (sasogeek):

x is in radians

OpenStudy (amistre64):

int by parts im assuming

OpenStudy (amistre64):

since sec^2 is the derivative of tan

OpenStudy (amistre64):

yeah; it ints up to \(e^{tan(x)}\)

OpenStudy (amistre64):

\[e^{tan(1)-e^{tan(0)}}\] \[e^{tan(1)}-1\]

OpenStudy (amistre64):

\[\int_{}Du.e^u.du \implies e^u\] \[u = tan(x) \implies du = sec^2(x)\] or some such logic right?

OpenStudy (angela210793):

Amistre will you please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please,please see my last comment please..I'd really appreciate it...:)

OpenStudy (amistre64):

lol..... sure

OpenStudy (anonymous):

u can also do a simple subsitution where u = tanx , therefore du = sec^2 x then plug in x=0 and x=1 to find the new values of u and integrate from 0 to pi/4 \[\int\limits_{0}^{\pi/4} e ^{u}du\]

OpenStudy (angela210793):

^_^ Thanks a looot! :)

OpenStudy (sasogeek):

I'm not getting it quite straight with the different explanations but i think I understand the logic behind each method... thanks :)

OpenStudy (amistre64):

yeah, the trick is to notice that the trig terms are really just derivatives of each other. this is really just asking you to find something like: \(\int_{}2x\) \(e^{x^2}\) dx

OpenStudy (anonymous):

this is a straight forward u - sub yes? \[u=tan(x)\] \[du=sec^2(x)dx\] but you should also change the limits of integration

OpenStudy (anonymous):

yes what kanade said exactly

OpenStudy (amistre64):

not even a u-sub; just a reintegration

OpenStudy (anonymous):

sorry i didn't see it

OpenStudy (sasogeek):

thanks :) much appreciated